OpenStudy (anonymous):

Describe the nature of the roots of the equation 8x^2-2x-1=0. a) 1 real rational root. b) 2 real irrational roots. c) 2 real rational roots. d) 2 imaginary conjugates.

OpenStudy (radar):

I would solve (or try to) and examine the roots.

OpenStudy (radar):

\[8x ^{2}-2x-1=0\]First is would attempt factoring (as it is easier for me than the quadratic formula)

OpenStudy (radar):

Yes, !! It does factor. (4x+1)(2x-1)=0 4x=-1, then x = -1/4 a real rational number (answer) other solution 2x-1=0 2x=1 x=1/2 another real rational number (answer)

OpenStudy (radar):

I realize the question was closed, but I believe this was your first question, and don't want you to become discouraged with open study. Welcome and good luck.